September 9, 2015 Calibration

From GlueXWiki
Jump to: navigation, search

GlueX Calibration Meeting
Wednesday, September 9, 2015
11:00 am, EDT
JLab: CEBAF Center, F326

Communication Information

Remote Connection

You can connect using BlueJeans using the meeting number 630 804 895 .       (Click "Expand" to the right for more details -->):

  1. Make sure you have created a BlueJeans account via your JLab CUE account using this link:

  2. Meeting ID: 630804895
    • (you may need to type this in, depending how you connect)

  3. If connecting via Web Browser: click this link (no passcode is needed):

  4. If connecting via iOS or Android App:
    • Use your JLab e-mail address to log in and then enter the meeting ID given above to join the meeting

  5. If connecting via Phone: Dial one of the following numbers and then enter the meeting ID above and hit "#" or "##"

  6. If connecting via Polycom unit:
    • Dial or
    • Enter meeting ID above
    • Use *4 to unmute


Talks can be deposited in the directory /group/halld/www/halldweb/html/talks/2015 on the JLab CUE. This directory is accessible from the web at .


  1. Announcements
  2. Summary of Detector Performance
  3. Subdetector Reports
  4. Simulations
  5. Calibration Processing
  6. AOB


Attending: Sean (NU); Simon, Luke, Alex B., Adesh, Eugene, Justin, Nathan, Beni, Paul, Fernando (JLab); Curtis, Will M, Naomi (CMU); Eric, Mahmoud (FIU); Brad (FSU)


  • Eugene had previously sent out an email detailing some changes being planned to the magnet, to improve the cooling performance. These changes won't be done until early December, then the magnet takes ~3 weeks to cool down, and 2-3 weeks will be needed for testing. So, the soonest one can expect the magnetic to be on and ready is mid-January. A meeting with local experts is scheduled for tomorrow to review these changes.
  • Curtis brought up the Fall Collaboration Meeting, which is in one month. Sean will give an overview of the status of calibrations, with more details being given in individual detector reports. It would be useful to know how much data is needed to complete the calibrations. Sean will also give an update of semi-automated calibration procedures.
  • Sean talked to Kei, who reported that the last offline monitoring run has been plagued with various problems, and about half of the jobs are left after almost two weeks. They agreed to scrap the current run and start a fresh one when Kei is back (this Friday, perhaps).

Summary of Detector Performance

Curtis has received updates for most detector systems, and is still waiting on updates from the beamline and online groups. Sean is looking more closely at the numbers to identify areas to focus on (e.g. efficiencies).

Subdetector Reports

  • FCAL - Adesh has been working mainly on hardware issues lately. No updates from other IU people either, as of last Friday.
  • BCAL - No one at the meeting, therefore no news.
  • CDC' - Beni mentioned that a lot of cosmic data was taken over the long Labor Day weekend, about 300M triggers in mode 7 with BCAL trigger. Will has started looking at this data for gain calibrations. Beni reported that Lubomir tried taking data with the CDC using an iron source inserted in the bore. He saw clear signals which (1) were smaller than expected, and (2) had reflections 200 ns later with 20% smaller peak. There is an elog entry describing this. Simon reported that the CDC sag saga continues - he has done detailed doca studies in 21 out of 28 rings, and hopes to report more results at the drift chamber meeting tomorrow.
  • Tracking - Simon changed the FCAL track/shower matching so that instead of matching to the shower position (energy-weighted average), he checks matches to individual blocks, which provides better matching to showers created by charged particles. A similar change had been made to the BCAL track/shower matching. This was motivated by noticing that showers from charged particles were often very non-symmetric.
  • TOF - Brad has successfully run the calibration codes, and Beni verified that he got consistent results, so he is ready to use this code in the future.
  • Start Counter - No news either.
  • TAGH - Nathan is working on updating the time offsets with respect to the RF time. He finds that using PS triggered data yields a clean sample for calibrations. He has also started to look at timewalk calibrations.
  • TAGM - Alex is using PSC triggered data to fine-tune the timewalk calibrations. Note that using the RF time as a reference he obtained a resolution of 120 ps. The useful Spring data was taken before he updated the TAGM gains, so the timewalk corrections for some columns need more data.
  • RF time - There is an outstanding problem with a significant difference in RF time reported between different crates (~8 ps). Fernando reported on some detailed measurements done with the highest precision oscilloscope at JLab. They find an intrinsic jitter of ~600 fs in the RF signal, a resolution of ~<5 ps in the MO (which prescales the RF signal), and ~<4 ps jitter in the clock FPGA algorithm. The time resolution obtained from TDCs is consistent with the intrinsic resolution of the chip. So, the inter-crate difference is still not understood. A GlueX note is in preparation.


Mark I. has tagged a new sim-recon version for the simulations requested by the BCAL group. Sean will work on preparing conditions for studies of future running. Eugene suggested that we explore photon flux intensities of up to the GlueX low intensity plan (10^7 photon/s in the coherent peak).